thabratt:

the slope of the line below is -5/7. write a point-slope equation of the line using the coordinates of the labeled point.

AZ:

Hello! What is the labeled point?

AZ:

To begin, the point-slope equation is \( y- y_1 = m(x -x_1)\) where m is the slope and in this question is \(-\dfrac{5}{7}\) and your point is \( (x_1, y_1)\)